The measure of one acute angle in a right triangle is four times the measure of other acute angle. Write and solve a system of equations to find the measures of the acute angles.
Answer by fcabanski(874)

Call the smaller acute angle x and the larger y


y=4x


The sum of the interior angles in a triangle is 180 so


x+y+90=180 and thus x+y=90.


Solve by substitution.


y=4x so


x+4x=90
5x=90
x=18


y=4x=4(18)=72

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com